OpenStudy (anonymous):

2/5 of 45 ? i will give medal

OpenStudy (anonymous):

45 x .4 in other words 40% of 45

OpenStudy (anonymous):

\[\frac{2}{5}\times 45\] cancel first multiply last and forget the decimals and percents

OpenStudy (anonymous):

its 18

OpenStudy (anonymous):

wow XD

OpenStudy (anonymous):

or just find out wha t45 is as a fraction and then multiply them

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the answer is 18. divide 45 by 5 then multiply that by 2

OpenStudy (anonymous):

\[\frac{2}{5}\times 45=\frac{2}{5}\times 5\times 9=2\times 9=18\]

OpenStudy (anonymous):

thank you @PilarS.xoxo

OpenStudy (anonymous):

never thought of it that way @PilarS.xoxo

OpenStudy (anonymous):

What is the sum of one over six plus one over three in simplest form? one over nine one over two four over seven two over nine

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

thx j bird

OpenStudy (anonymous):

anyone know anything about axioms??

OpenStudy (anonymous):

otf!

OpenStudy (anonymous):

what does that mean??